# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 1 Introduction
Although these estimators have the same asymptotic normal distribution, their performance in moderate samples has been observed to vary in simulations. For example, cross-fitting, a general technique that is similar to sample splitting and applicable to many nonparametric estimators involving nuisance function estimation, has been empirically found to improve the performance in moderate samples (e.g., Li et al.,, 2022; Smith et al.,, 2024). Currently, the most widely accepted explanation is that cross-fitting removes a Donsker condition (van der Vaart and Wellner,, 1996; Kosorok,, 2008), which restricts the flexibility of the nuisance function estimator and is required when not adopting sample-splitting to control an empirical process term. However, if the Donsker condition is verified, it is theoretically unclear whether cross-fitting improves performance.
@@
In this paper, I attempt to compare the moderate-sample performance of AIPW estimators with and without cross-fitting from another perspective on the asymptotic behavior, the convergence of the corresponding Wald-CI coverage to nominal coverage, namely a Berry-Esseen-type bound, using an estimated asymptotic variance. This may be among the most important theoretical questions regarding these flexible nonparametric procedures following asymptotic normality, because one of the most common use of these nonparametric estimators in practice is the resulting asymptotically valid inference based on the corresponding Wald-CI. A few existing works provide Berry-Esseen-type bounds for nonparametric estimators. Quintas-Martinez, (2022); Chernozhukov et al., (2023) derived Berry-Esseen-type bounds for debiased machine learning. Because the true asymptotic variance was used to standardize the estimator, these results characterize the convergence rate of the estimator’s sampling distribution to the normal distribution, but do not directly characterize the confidence interval coverage when the asymptotic variance is estimated. Saco, (2025) characterized the convergence of confidence interval coverage to nominal coverage for debiased machine learning in the special case of partially linear models. For relatively simpler procedures, Berry-Esseen-type bounds (or more precise characterizations in the form of Edgeworth expansions) have been established. Hall, (1992) reviewed these results for the sample mean, some parametric models, and kernel smoothing. Berry-Eseen-type bounds for U-statistics (e.g., Callaert et al.,, 2007; Bickel et al.,, 2007), the sample mean of high-dimensional random vectors (e.g., Zhilova,, 2020, 2022), stationary dependent data (e.g., Bentkus et al.,, 1997; Jirak,, 2016, 2023), or design based average treatment effect (ATE) estimation (e.g., Shi and Ding,, 2022), have also been derived. However, to the best of my knowledge at the time of writing, no existing literature addresses the convergence rate of CI coverage for a flexible nonparametric ATE estimator with an estimated variance.
@@
I consider a simple but practically relevant toy example, the estimation of the ATE with an AIPW estimator, with or without cross-fitting, in a randomized controlled trial (RCT). Generally, the AIPW estimator involves two important nuisance functions, the outcome model and the propensity score. In contrast to the observational data setting, in the RCT setting, although the outcome model may be estimated almost arbitrarily, the propensity score is known and does not need to be estimated or calibrated. Because of the known propensity score, the RCT setting completely differs from the observational setting, the focus of most above theoretical studies. There is no issue of the non-existence of root-$n$ consistent estimators when both nuisance functions are unknown in CODA asymptotic theory (Theorem 2 of Robins and Ritov,, 1997). The focus of Liu et al., (2020, 2023), a potentially problematic bias term or its Cauchy-Schwarz upper bound, also equals zero. I consider a plug-in estimator of the asymptotic variance when computing a standard error to construct a Wald-CI. This is the most common and widely applicable standard error in the literature. Efficient asymptotic variance estimators in Tran et al., (2023); Van der Laan and Rose, (2018) only apply to binary outcomes. For general outcomes, these estimators may need additional nuisance estimators, and thus are less computationally convenient and not studied in this paper. Although this example appears simple and thoroughly studied, I show new and non-trivial findings regarding Wald-CI coverage.
@@
    I derive the first non-asymptotic model-agnostic Berry-Esseen-type bounds for AIPW estimators in RCTs when the variance is estimated (Theorems S1 and S3 in Section S2 of the Supplementary Material). For better readability, I present the asymptotic version in Theorems 1 and 3 in the main text. The bound mainly depends on whether cross-fitting is adopted, the convergence behavior of the outcome model estimator, and the complexity of the Donsker class when cross-fitting is not adopted. The best-case rate is ${\mathrm{O}}(n^{-1/2})$, same as the Berry-Esseen Theorem, up to a log factor. Although the known propensity score resolves many issues studied in existing literature, my Berry-Esseen-type bounds still suggest potential differences between AIPW estimators with and without cross-fitting in terms of Wald-CI coverage due to different bounds on empirical processes. Section 3.3 contains a discussion of the interpretation of these bounds.
@@
    I derive non-asymptotic model-agnostic bounds on the difference between (i) the expectation of the plug-in variance estimators, and (ii) an oracle variance of the AIPW estimator based on a finite-sample deterministic approximation to the outcome estimator, with and without cross-fitting (Theorems S2 and S4 in Section S2 of the Supplementary Material). Again, I present asymptotic versions Theorem 2 and 4 in the main text. Somewhat surprisingly, with a known propensity score and for a reasonably chosen oracle variance, the difference vanishes at a rate that is almost second order, rather than first order, regarding the $L_{2}$-distance between the outcome model estimator and deterministic approximation. That is, the bias of the plug-in variance estimator for the oracle variance vanishes relatively fast in RCTs.
@@
  5. 5.
    With the above results, I provide new theoretical justification for adopting cross-fitting when using AIPW estimators in RCTs, even if Donsker conditions are verified.
@@ 2 Preliminaries
Suppose that a prototypical full data point consists of a vector of baseline covariates $X\in\mathcal{X}$, a binary treatment $A\in\{0,1\}$, and real-valued counterfactual outcomes $Y(a)$ corresponding to treatment status $a\in\{0,1\}$ (Neyman,, 1923; Rubin,, 1974). A prototypical observed data point consists of $V:=(X,A,Y)$ where $Y=Y(A)$. Suppose that the data consists of $n$ independent and identically distributed (i.i.d.) copies $(V_{i})_{i=1}^{n}$ of $V$ drawn from a true distribution $P_{*}$. I consider an RCT setting where treatment $A$ is randomized according to $\Pr(A=a\mid X=x)=\pi_{*}(a\mid x)$ for a known propensity score function $\pi_{*}$. The most common choice of $\pi_{*}$ is a constant function.
@@
Let $Q_{*,a}:x\mapsto{\mathbb{E}}[Y\mid X=x,A=a]$ be the true outcome model and $\psi_{*,a}:={\mathbb{E}}[Y(a)]$ be the mean counterfactual outcome corresponding to treatment $a$, which equals ${\mathbb{E}}[Q_{*,a}(X)]$ (Robins,, 1986). These estimands are building blocks for common treatment effect estimates, including risk difference $\psi_{*,1}-\psi_{*,0}$, risk ratio $\psi_{*,1}/\psi_{*,0}$ (when $Y\geq 0$), and odds ratio $\{\psi_{*,1}/(1-\psi_{*,1})\}/\{\psi_{*,0}/(1-\psi_{*,0})\}$ (when $Y\in[0,1]$).
@@
I will adopt empirical process notations (van der Vaart and Wellner,, 1996): For any distribution $P$ and $P$-integrable function $f$, let $Pf:=\int f{\mathrm{d}}P$. The empirical distribution is denoted by $P_{n}$. If the data is split into $K$ (disjoint) folds, I use $I_{k}$ to denote the index of the data points in fold $k$ so that $\sqcup_{k=1}^{K}I_{k}=\{1,\ldots,n\}$, and use $P_{n,k}$ to denote the empirical distribution in fold $k$. For any probability measure $P$ and $q\geq 1$, let $\|\cdot\|_{q,P}$ denote the $L_{q}(P)$-norm of a function, that is, $\|f\|_{q,P}:=(P|f|^{q})^{1/q}$, and let $\|\cdot\|_{q}:=\|\cdot\|_{q,P_{*}}$.
@@
All asymptotics refer to the scenario where the sample size $n\to\infty$ and all other aspects of the data-generating mechanism are fixed. I may use $n\to\infty$ to denote this asymptotic scenario for short. I adopt big-O and little-o notations in asymptotic scenarios, and use $\Theta$ to denote the same order; that is, for any real sequences $a_{n}$ and $b_{n}$, $a_{n}=\Theta(b_{n})$ if $a_{n}={\mathrm{O}}(b_{n})$ and $b_{n}={\mathrm{O}}(a_{n})$. I write $a_{n}=\Theta_{+}(b_{n})$ if $a_{n}=\Theta(b_{n})$ and $a_{n}\geq 0$. If cross-fitting is adopted, to study the impact of the choice of $K$, the number of folds $K$ is not viewed as an absolute constant.
@@
An estimator $\hat{\theta}$ of $\theta$ is said to be asymptotically linear with influence function $f$ if $P_{*}f=0$, $P_{*}f^{2}<\infty$, and $\hat{\theta}=\theta+P_{n}f+{\mathrm{o}}_{p}(n^{-1/2})$. For any $\alpha\in(0,1)$, let $z_{\alpha}$ denote the $(1-\alpha)$-quantile of standard Gaussian. Let $\phi$ denote the standard Gaussian density. I use $\mathcal{C}$} to denote an absolute positive constant that may vary line by line. For a random function $\hat{f}$, I use $^{\mathbb{E}}_{\hat{f}}$ to denote the expectation over $\hat{f}$ only. I use $\mathbbm{1}$} to denote the indicator function.
@@
A non-cross-fit AIPW estimator $\tilde{\psi}_{a}$ (Robins et al.,, 1994, 1995) of $\psi_{*,a}$ is constructed as $P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})$, where, for any function $Q:\mathcal{X}\to{\mathbb{R}}$ and $a^{\prime}\in\{0,1\}$, I define transformation ${\mathcal{T}}_{a^{\prime}}(Q):v=(x,a,y)\mapsto{\mathbbm{1}}(a=a^{\prime})\{y-Q(x)\}/\pi_{*}(a^{\prime}\mid x)+Q(x)$, and $\tilde{Q}_{a}$ is a (possibly black-box) flexible estimator of $Q_{*,a}$. Under somewhat mild conditions below, $\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})$ converges to a normal distribution. Usually, a consistent estimator of the asymptotic variance is computed based on the estimated influence function, namely $\tilde{\sigma}_{a}^{2}:=P_{n}({\mathcal{T}}_{a}(\tilde{Q}_{a})-\tilde{\psi}_{a})^{2}$, so that $\tilde{\sigma}_{a}/\sqrt{n}$ is asymptotically valid standard error to construct a Wald-CI. For example, $\tilde{\psi}_{a}\pm z_{\alpha/2}\tilde{\sigma}_{a}/\sqrt{n}$ is the $(1-\alpha)$-level symmetric two-sided Wald-CI.
@@
Motivated by such robustness against arbitrary bias in $\tilde{Q}_{a}$ and $\hat{Q}_{k,a}$, for each $a\in\{0,1\}$, I consider a fixed function $Q_{\#,a}:\mathcal{X}\to{\mathbb{R}}$ in $L_{2}(P_{*})$ that may depend on the sample size $n$ and may differ from $Q_{*,a}$ and $Q_{\infty,a}$. Since I analyze estimators with and without cross-fitting separately, $Q_{\#,a}$ may differ in the two cases. The theoretical results below hold for any fixed $Q_{\#,a}$ sufficiently close to $\tilde{Q}_{a}$ and $\hat{Q}_{k,a}$, but the bounds are the tightest when $Q_{\#,a}$ is a good non-asymptotic approximation to $\tilde{Q}_{a}$ or $\hat{Q}_{k,a}$. In particular, $Q_{\#,a}$ may be taken as $x\mapsto{\mathbb{E}}[\tilde{Q}_{a}(x)]$ for the non-cross-fit estimator, and $x\mapsto{\mathbb{E}}[\hat{Q}_{k,a}(x)]$ for the cross-fit estimator if $\hat{Q}_{k,a}$ ($k=1,\ldots,K$) has a common mean. This setup enables a sample-size-dependent deterministic approximation to the outcome model estimator, highlighting the non-asymptotic nature of the bounds below. As shown in the simulation in Section 5, a sample-size-dependent approximation can be much more accurate than the asymptotic approximation. I assume that ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ or ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ ($k=1,\ldots,K$), depending on the estimator being considered, is non-zero, namely the outcome model estimator is non-degenerate. Otherwise, the AIPW estimators reduce to trivial sample means of i.i.d. random variables, and results on Wald-CI coverage have been established with Edgeworth expansion (Hall,, 1992; Bhattacharya and Ghosh,, 2007). In the asymptotic scenario $n\to\infty$, I assume that ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ or ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ ($k=1,\ldots,K$) converges to zero.
@@
$\sigma_{\dagger,a}^{2}:=\begin{cases}{\mathbb{E}}[\tilde{\sigma}_{a}^{2}]&\text{for non-cross-fit AIPW estimator}\\ {\mathbb{E}}[\hat{\sigma}_{a}^{2}]&\text{for cross-fit AIPW estimator}\end{cases}$
@@
The positivity in Conditions 1 and 5 holds if $\Pr({\mathrm{Var}}(Y\mid X,A=a)>0)>0$. Under Condition 5, with probability one, $P_{*}D_{a}(\tilde{Q}_{a},\psi_{*,a})^{2}$ and $P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}$ ($k=1,\ldots,K$) are bounded, say by $\bar{\sigma}_{a}^{2}$; $P_{*}|D_{a}(\tilde{Q}_{a},\psi_{*,a})|^{3}$ and $P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}$ ($k=1,\ldots,K$) are bounded, say by $\bar{\rho}_{a}$. The boundedness in Conditions 2, 4 and 5 holds (or is anticipated to hold) if the outcome $Y$ is bounded. Condition 3 holds in almost all RCTs.

### v2

@@ 1 Introduction
Although these estimators have the same asymptotic normal distribution, their performance in moderate samples has been observed to vary in simulations. For example, cross-fitting, a general technique that is similar to sample splitting and applicable to many nonparametric estimators involving nuisance function estimation, has been empirically found to improve the performance in moderate samples (e.g., Li et al.,, 2022; Smith et al.,, 2024). Currently, the most widely accepted explanation is that cross-fitting removes a Donsker condition (Kosorok,, 2008; van der Vaart and Wellner,, 2023), which restricts the flexibility of the nuisance function estimator and is required when not adopting sample-splitting to control an empirical process term. However, if the Donsker condition is verified, it is theoretically unclear whether cross-fitting improves performance.
@@
In this paper, I attempt to compare the moderate-sample performance of AIPW estimators with and without cross-fitting from another perspective on the asymptotic behavior, the convergence of the corresponding Wald-CI coverage to nominal coverage, namely a Berry-Esseen-type bound, using an estimated asymptotic variance. This may be among the most important theoretical questions regarding these flexible nonparametric procedures following asymptotic normality, because one of the most common use of these nonparametric estimators in practice is the resulting asymptotically valid inference based on the corresponding Wald-CI. A few existing works provide Berry-Esseen-type bounds for nonparametric estimators. Quintas-Martinez, (2022); Chernozhukov et al., (2023) derived Berry-Esseen-type bounds for debiased machine learning. Because the true asymptotic variance was used to standardize the estimator, these results characterize the convergence rate of the estimator’s sampling distribution to the normal distribution, but do not directly characterize the CI coverage when the asymptotic variance is estimated. Saco, (2025) characterized the convergence of CI coverage to nominal coverage for debiased machine learning in the special case of partially linear models. For relatively simpler procedures, Berry-Esseen-type bounds on CI coverage rates (or more precise characterizations of the rates) have been established. Hall, (1992) reviewed these results for the sample mean, some parametric models, and kernel smoothing. Berry-Eseen-type bounds for U-statistics (e.g., Callaert et al.,, 2007; Bickel et al.,, 2007), the sample mean of high-dimensional random vectors (e.g., Zhilova,, 2020, 2022), stationary dependent data (e.g., Bentkus et al.,, 1997; Jirak,, 2016, 2023), or design based average treatment effect (ATE) estimation (e.g., Shi and Ding,, 2022), have also been derived. However, to the best of my knowledge at the time of writing, no existing literature addresses the convergence rate of CI coverage for a flexible nonparametric ATE estimator with an estimated variance.
@@
I consider a simple but practically relevant example, the estimation of the ATE with an AIPW estimator, with or without cross-fitting, in a randomized controlled trial (RCT). Generally, the AIPW estimator involves two important nuisance functions, the outcome model and the propensity score. In contrast to the observational data setting, in the RCT setting, the propensity score is known and need not be estimated or calibrated. Because of this knowledge, the AIPW estimator behaves fundamentally differently in the RCT setting. It is fully robust against inconsistent estimation of the outcome model and may achieve efficiency gains, making it a potentially desirable estimator in RCTs. Moreover, there is no issue of the non-existence of root-$n$ consistent estimators when both nuisance functions are unknown in CODA asymptotic theory (Theorem 2 in Robins and Ritov,, 1997). The focus of Liu et al., (2020, 2023), a potentially problematic bias term or its Cauchy-Schwarz upper bound, also equals zero. Because of the additional substantive complexities in the observational setting, I confine attention to RCTs.
I consider a plug-in estimator of the asymptotic variance when computing a standard error to construct a Wald-CI. This is the most common and widely applicable standard error in the literature. Efficient asymptotic variance estimators in Tran et al., (2023); Van der Laan and Rose, (2018) only apply to binary outcomes. For general outcomes, these estimators may need additional nuisance estimators, and thus are less computationally convenient and not studied in this paper. Although this example appears simple and thoroughly studied, I show new and non-trivial findings regarding Wald-CI coverage.
@@
    I derive the first non-asymptotic model-agnostic Berry-Esseen-type bounds for AIPW estimators with black-box nuisance estimators in RCTs when the variance is estimated (Theorems S1 and S3 in Section S1 of the Supplementary Material). For better readability, I present the asymptotic versions in Theorems 1 and 3 in the main text. The rate depends on whether cross-fitting is adopted, the convergence behavior of the outcome model estimator, and the complexity of the Donsker class when cross-fitting is not adopted. The best-case rate is root-$n$ up to a log factor, which is achieved under sub-Weibull conditions and when the outcome model estimator converges to some function at the root-$n$-rate. This rate is comparable to the Berry-Esseen Theorem for the sample mean. Although the known propensity score resolves many issues studied in existing literature, my Berry-Esseen-type bounds still suggest potential differences between AIPW estimators with and without cross-fitting in terms of Wald-CI coverage due to different bounds on empirical processes. Section 3.3 contains a discussion of the interpretation of these bounds.
@@
    I derive non-asymptotic model-agnostic bounds on the difference between (i) the expectation of the plug-in variance estimators, and (ii) an oracle variance of the AIPW estimator based on a finite-sample deterministic approximation to the outcome estimator, with and without cross-fitting (Theorems S2 and S4 in Section S1 of the Supplementary Material). I present asymptotic versions in Theorems 2 and 4 in the main text. Somewhat surprisingly, with a known propensity score and for a reasonably chosen oracle variance, the difference vanishes at a rate that is almost second order, rather than first order, regarding the $L_{2}$-distance between the outcome model estimator and deterministic approximation. That is, the bias of the plug-in variance estimator for the oracle variance vanishes relatively fast in RCTs.
@@
With the above results, I provide new theoretical justification for adopting cross-fitting when using AIPW estimators in RCTs, even if Donsker conditions hold.
@@ 2 Preliminaries
Suppose that a prototypical full data point consists of a vector of baseline covariates $X\in\mathcal{X}$, a binary treatment $A\in\{0,1\}$, and real-valued counterfactual outcomes $Y(a)$ corresponding to treatment status $a\in\{0,1\}$ (Neyman,, 1923; Rubin,, 1974). A prototypical observed data point consists of $V:=(X,A,Y)$ where $Y=Y(A)$. Suppose that the data consists of $n$ independent and identically distributed (i.i.d.) copies $(V_{i})_{i=1}^{n}$ of $V$ drawn from a true distribution $P_{*}$. I consider an RCT setting where treatment $A$ is randomized according to ${\mathrm{pr}}(A=a\mid X=x)=\pi_{*}(a\mid x)$ for a known propensity score function $\pi_{*}$. The most common choice of $\pi_{*}$ is a constant function.
@@
Let $Q_{*,a}:x\mapsto{E}[Y\mid X=x,A=a]$ be the true outcome model and $\psi_{*,a}:={E}[Y(a)]$ be the mean counterfactual outcome corresponding to treatment $a$, which equals ${E}[Q_{*,a}(X)]$ (Robins,, 1986). These estimands are building blocks for common treatment effect estimates, including risk difference $\psi_{*,1}-\psi_{*,0}$, risk ratio $\psi_{*,1}/\psi_{*,0}$ (when $Y\geq 0$), and odds ratio $\{\psi_{*,1}/(1-\psi_{*,1})\}/\{\psi_{*,0}/(1-\psi_{*,0})\}$ (when $Y\in[0,1]$).
@@
I will adopt empirical process notations (van der Vaart and Wellner,, 2023): For any distribution $P$ and $P$-integrable function $f$, let $Pf:=\int f{\mathrm{d}}P$. The empirical distribution is denoted by $P_{n}$. If the data is split into $K$ (disjoint) folds, I use $I_{k}$ to denote the index of the data points in fold $k$ so that $\sqcup_{k=1}^{K}I_{k}=\{1,\ldots,n\}$, and use $P_{n,k}$ to denote the empirical distribution in fold $k$. For any probability measure $P$ and $q\geq 1$, let $\|\cdot\|_{q,P}$ denote the $L_{q}(P)$-norm of a function, that is, $\|f\|_{q,P}:=(P|f|^{q})^{1/q}$, and let $\|\cdot\|_{q}:=\|\cdot\|_{q,P_{*}}$.
@@
All asymptotics refer to the scenario where the sample size $n\to\infty$ and all other aspects of the data-generating mechanism are fixed. I may use $n\to\infty$ to denote this asymptotic scenario for short. I adopt big-O and little-o notations in asymptotic scenarios, and use $\Theta$ to denote the same order; that is, for any real sequences $a_{n}$ and $b_{n}$, $a_{n}=\Theta(b_{n})$ if $a_{n}={\mathrm{O}}(b_{n})$ and $b_{n}={\mathrm{O}}(a_{n})$. I write $a_{n}=\Theta_{+}(b_{n})$ if $a_{n}=\Theta(b_{n})$ and $a_{n}\geq 0$. If cross-fitting is adopted, the number of folds $K$ may vary with $n$.
@@
An estimator $\hat{\theta}$ of $\theta$ is said to be asymptotically linear with influence function $f$ if $P_{*}f=0$, $P_{*}f^{2}<\infty$, and $\hat{\theta}=\theta+P_{n}f+{\mathrm{o}}_{p}(n^{-1/2})$. For any $\alpha\in(0,1)$, let $z_{\alpha}$ denote the $(1-\alpha)$-quantile of standard Gaussian. Let $\phi$ denote the standard Gaussian density. I use $\mathcal{C}$} to denote an absolute positive constant that may vary line by line. For a random function $\hat{f}$, I use $^{E}_{\hat{f}}$ to denote the expectation over $\hat{f}$ only. I use $\mathbbm{1}$} to denote the indicator function.
@@
A non-cross-fit AIPW estimator $\tilde{\psi}_{a}$ (Robins et al.,, 1994, 1995) of $\psi_{*,a}$ is constructed as $P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})$, where, for any function $Q:\mathcal{X}\to{\mathbb{R}}$ and $a^{\prime}\in\{0,1\}$, I define transformation ${\mathcal{T}}_{a^{\prime}}(Q):v=(x,a,y)\mapsto{\mathbbm{1}}(a=a^{\prime})\{y-Q(x)\}/\pi_{*}(a^{\prime}\mid x)+Q(x)$, and $\tilde{Q}_{a}$ is a (possibly black-box) flexible estimator of $Q_{*,a}$. Under somewhat mild conditions below, $n^{1/2}(\tilde{\psi}_{a}-\psi_{*,a})$ converges to a normal distribution. Usually, a consistent estimator of the asymptotic variance is computed based on the estimated influence function, namely $\tilde{\sigma}_{a}^{2}:=P_{n}({\mathcal{T}}_{a}(\tilde{Q}_{a})-\tilde{\psi}_{a})^{2}$, so that $\tilde{\sigma}_{a}/\sqrt{n}$ is asymptotically valid standard error to construct a Wald-CI. For example, $\tilde{\psi}_{a}\pm n^{-1/2}z_{\alpha/2}\tilde{\sigma}_{a}$ is the $(1-\alpha)$-level symmetric two-sided Wald-CI.
@@
Motivated by such robustness against arbitrary bias in $\tilde{Q}_{a}$ and $\hat{Q}_{k,a}$, for each $a\in\{0,1\}$, I consider a fixed function $Q_{\#,a}:\mathcal{X}\to{\mathbb{R}}$ in $L_{2}(P_{*})$ that may depend on the sample size $n$ and may differ from $Q_{*,a}$ and $Q_{\infty,a}$. Since I analyze estimators with and without cross-fitting separately, $Q_{\#,a}$ may differ in the two cases. The theoretical results below hold for any fixed $Q_{\#,a}$ sufficiently close to $\tilde{Q}_{a}$ and $\hat{Q}_{k,a}$, but the bounds are the tightest when $Q_{\#,a}$ is a good non-asymptotic approximation to $\tilde{Q}_{a}$ or $\hat{Q}_{k,a}$. In particular, $Q_{\#,a}$ may be taken as $x\mapsto{E}[\tilde{Q}_{a}(x)]$ for the non-cross-fit estimator, and $x\mapsto{E}[\hat{Q}_{k,a}(x)]$ for the cross-fit estimator if $\hat{Q}_{k,a}$ ($k=1,\ldots,K$) has a common mean. This setup enables a sample-size-dependent deterministic approximation to the outcome model estimator, highlighting the non-asymptotic nature of the bounds below. As shown in the simulation in Section 5, a sample-size-dependent approximation can be much more accurate than the asymptotic approximation. I assume that ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ or ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ ($k=1,\ldots,K$), depending on the estimator being considered, is non-zero, namely the outcome model estimator is non-degenerate. Otherwise, the AIPW estimators reduce to trivial sample means of i.i.d. random variables, and results on Wald-CI coverage have been established with Edgeworth expansion (Hall,, 1992; Bhattacharya and Ghosh,, 2007). In the asymptotic scenario $n\to\infty$, I assume that ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ or ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ ($k=1,\ldots,K$) converges to zero.
@@
$\sigma_{\dagger,a}^{2}:=\begin{cases}{E}[\tilde{\sigma}_{a}^{2}]&\text{for non-cross-fit AIPW estimator}\\ {E}[\hat{\sigma}_{a}^{2}]&\text{for cross-fit AIPW estimator}\end{cases}$
@@
The positivity in Conditions 1 and 5 holds if ${\mathrm{pr}}({\mathrm{var}}(Y\mid X,A=a)>0)>0$. Under Condition 5, with probability one, $P_{*}D_{a}(\tilde{Q}_{a},\psi_{*,a})^{2}$ and $P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}$ ($k=1,\ldots,K$) are bounded, say by $\bar{\sigma}_{a}^{2}$; $P_{*}|D_{a}(\tilde{Q}_{a},\psi_{*,a})|^{3}$ and $P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}$ ($k=1,\ldots,K$) are bounded, say by $\bar{\rho}_{a}$. The boundedness in Conditions 2, 4 and 5 holds (or is anticipated to hold) if the outcome $Y$ is bounded. Condition 3 holds in almost all RCTs.